OpenStudy (anonymous):

whats the correct equation (in slope-intercept form). 2x + 3y = 6

OpenStudy (anonymous):

2x + 3y = 6 So rearrange for y=mx + c 2x + 3y = 6 3y = -2x + 6 Y= -2/3x + 2 The gradient/slope is in front of x which is -2/3

OpenStudy (anonymous):

Slope-intercept form is \[y = mx + b\] Where m is the slope, and b is where the line intercepts the y axis. Use algebra to rearrange the original equation so that it conforms to this format.\[3y=-2x+6\]\[y = -\frac{2}{3}x+2\]So the slope is -2/3 and the y intercept is 2